The slope of the line which passes through the point (8,14) and (0,4) is


1
Expert's answer
2021-05-11T13:31:39-0400

Equation of line is "y=kx+c"

We know two points on this line, so we can make the system of equations:

"\begin{cases}
14=8k+c \\
4=c
\end{cases} =>
k=\frac{10}8=\frac{5}4"

Next, "\arctan(k) =\arctan(1.25)".The slope is "k=\frac{5}{4}."